OpenStudy (anonymous):

a 232lb fullback runs 40yd at 19.8 +/- 0.1mi/h (b) what is his uncertainty in his position?

OpenStudy (aaronq):

use: \(\Delta x \Delta p \ge h\) where, x= position, p=momentum = velocity*mass \(\Delta x \Delta mv \ge\dfrac{h}{4\pi} \rightarrow\Delta x \ge \dfrac{h}{4\pi mv} \)